# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ 3.2 The Cauchy class
$\rho_{h}=(1+|h|^{2\alpha+1})^{-\beta/(2\alpha+1)}$ (3.6)
@@
Motivated by its ability to decouple the short- and long-term persistence, this process was studied as a empirical model for log-spot variance in Bennedsen, Lunde, and Pakkanen (2022). They fitted it with method of moments to high-frequency data from the E-mini S&P 500 futures contract and indeed found evidence of both roughness and long memory.
4 Monte Carlo simulation
@@
The fOU model was studied in Bolko, Christensen, Pakkanen, and Veliyev (2023) and Wang, Xiao, and Yu (2023) to describe the stochastic log-variance of financial asset returns, $Y_{t}=\log\sigma_{t}^{2}$. We follow this analogue here and adopt the Monte Carlo design of the former article by examining five different settings for $\theta_{\text{OU}}$ (listed in Table 1) ranging from very rough and short memory ($\alpha=-0.45$ or $H=0.05$) to smooth and long memory ($\alpha=0.20$ or $H=0.70$). We include an equal number of distinct parameter values for $\theta_{\text{CC}}$ (listed in Table 3), keeping the roughness index fixed at the values of the Hurst exponent from the fOU process, i.e. $\alpha=-0.45,-0.40,\dots,0.20$, while moving the long-run persistence in the opposite direction, i.e. $\beta=0.25,0.50,\dots,1.25$. This design ensures that the latter features multiple settings with both roughness and long memory.
We simulate over the interval $[0,T]$, where $T$ can be interpreted as the number of days, setting $T=1095,1825,2555$. This amounts to three, five, and seven years worth of data. The intermediate value roughly matches the sample size in our empirical application. We assume the process is observed discretely on an equispaced partition of $N$ points per day, so the $n=N\times T+1$ data vector is $(Y_{t-1+i\Delta})_{t=1,\dots,T\text{ and }i=0,1,\dots,N}$, where $\Delta=1/N$. We follow Christensen, Thyrsgaard, and Veliyev (2019) and Li, Todorov, and Tauchen (2013) and base the analysis on $N=12$ observations per unit interval, yielding a new observation of the log-spot variance every second hour.
We simulate the Cauchy class via circulant embedding (e.g., Asmussen and Glynn, 2007), which is an exact discretization that runs in $O(n\log n)$ time. However, since the acf of the fOU needs to be evaluated with the help of numerical integration we opt for a slightly different approach for that model. In particular, at a generic time $t$, the solution of the fOU process in (3.4) implies that:
@@
To implement CL, we need to decide the terms to include in (2.4) through the set $Q$. We experiment with triwise (i.e. $q^{j}=q=3$) CLE with $K=18$, where the elements of $k^{j}$ are separated by a fixed amount, i.e. $k^{j}=(0,\ell,2\ell)$ for $\ell=1,\dots,10$, 50, 100, 200, 500, 1000, 2000, 5000, 10000. This setup means that $cl(\theta;y)$ is a summation of densities of the form $f(y_{i},y_{i+\ell},y_{i+2\ell};\theta)$. The reported results are rather robust to the exact configuration of the CL procedure.
@@
We benchmark MCLE against a method-of-moments estimator (MME). We follow the implementation of Wang, Xiao, and Yu (2023) for the fOU process and Bennedsen, Lunde, and Pakkanen (2022) for the Cauchy class. The details are reviewed in Appendix B. It is worth pointing out that the first-step MME estimator of $\alpha$ (or $H$) relies on change-of-frequency approach based on first-differenced data. Hence, it does not depend on whether the mean is known or estimated. We should also emphasize that we start the MCLE routine with the MME as initial value.
In the left-hand side of Table 12, we report the Monte Carlo average (standard error in parenthesis) of each parameter estimate for the fOU process, while Table 34 presents the corresponding results for the Cauchy class. The MME is shown in the right-hand side of each table. The number of Monte Carlo replications is 10,000.
Table 1: Parameter estimation of the fOU process ($\mu$ known).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
$\mu$ 0.000 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000)
$\kappa$ 0.070 0.0823 (0.0319) 0.0726 (0.0180) 0.0718 (0.0138) 0.0732 (0.0157) 0.0720 (0.0122) 0.0715 (0.0103)
$\nu$ 0.200 0.2032 (0.0245) 0.2008 (0.0129) 0.2004 (0.0096) 0.2000 (0.0079) 0.2000 (0.0061) 0.2000 (0.0052)
$\alpha$ 0.200 0.2015 (0.0276) 0.1995 (0.0187) 0.1995 (0.0146) 0.1998 (0.0109) 0.1998 (0.0085) 0.1999 (0.0071)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval, corresponding to an observation every second hour. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 2: Parameter estimation of the fOU process ($\mu$ estimated).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
$\mu$ 0.000 0.0003 (0.0691) 0.0002 (0.0591) 0.0001 (0.0538) 0.0003 (0.0688) 0.0002 (0.0589) 0.0001 (0.0534)
$\kappa$ 0.070 0.0837 (0.0177) 0.0745 (0.0130) 0.0735 (0.0109) 0.0795 (0.0158) 0.0765 (0.0120) 0.0752 (0.0101)
$\nu$ 0.200 0.2007 (0.0073) 0.1999 (0.0058) 0.1999 (0.0049) 0.2000 (0.0079) 0.2000 (0.0061) 0.2000 (0.0052)
$\alpha$ 0.200 0.1937 (0.0146) 0.1945 (0.0123) 0.1955 (0.0101) 0.1998 (0.0109) 0.1998 (0.0085) 0.1999 (0.0071)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval, corresponding to an observation every second hour. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 3: Parameter estimation of the Cauchy class ($\mu$ known).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
$\mu$ 0.000 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000)
$\beta$ 1.250 1.3442 (0.1070) 1.2542 (0.0788) 1.2529 (0.0672) 1.2885 (0.1847) 1.2788 (0.1471) 1.2757 (0.1257)
$\nu$ 0.200 0.1921 (0.0056) 0.1999 (0.0046) 0.1999 (0.0039) 0.1999 (0.0058) 0.1999 (0.0045) 0.2000 (0.0039)
$\alpha$ 0.200 0.1984 (0.0085) 0.2000 (0.0065) 0.2000 (0.0055) 0.2406 (0.0109) 0.2407 (0.0085) 0.2408 (0.0072)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval, corresponding to an observation every second hour. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 4: Parameter estimation of the Cauchy class ($\mu$ estimated).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval, corresponding to an observation every second hour. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Looking at Table 12 first, we observe that both MCLE and MME perform well. The parameters are estimated with an immaterial bias, even with a limited sample. The sole exception is the mean-reversion parameter, $\kappa$, which is estimated with an upward bias, but it vanishes as the sample grows and is negligible for $T=2{,}555$. The standard errors are small and MCLE is generally more efficient than MME.
@@
Turning next to Table 34 for the Cauchy class, the conclusion is two-fold. With a known mean, the distribution of the covariance-related parameter estimators is again centered around the true value with minimal sampling variation for MCLE. The main difference is that we observe a slight upward bias in the $\beta$ parameter, although it is within a Monte Carlo standard error. By contrast, the MME of $\beta$ is severely downward biased and appears to move in the wrong direction in the presence of strong persistence ($\beta\ll 1$). Although this observation is present for all the rough settings, it is most pronounced for $\alpha=-0.45$ and $\alpha=-0.40$. The cause of this is a too low estimate of $\alpha$, which is input into the calculation of $\beta$. The first-stage MME of the roughness parameter relies on infill asymptotic theory, and it turns out to be severely biased for a fixed sample size. In turn, this distorts the estimates of the remaining parameters. This highlights a significant drawback of the two-step MME procedure. [^12] To investigate this issue further, in Appendix C we report simulation results from a Monte Carlo study, where the volatility process is sampled once per day, i.e. $N=1$. This shows that the bias problems in $\alpha$ with MME are intensified, indicating that it may be more imprecise with infrequently sampled data. This is especially troublesome, since the MME is often applied to daily realized variance (e.g. Gatheral, Jaisson, and Rosenbaum, 2018; Bennedsen, Lunde, and Pakkanen, 2022; Wang, Xiao, and Yu, 2023).
@@
However, when the mean is estimated, even more than before the MCLE exhibits a pronounced upward bias for $\beta$ in the long memory configurations, although it again dissipates as the sample size is increased. The estimator of the mean—the sample average—is unbiased across replica, but it exhibits a great deal of sampling variation and is typically way off target in each individual simulation, as evident from the magnitude of the standard error. In essence, the sample average is a noisy estimator of the mean for persistent processes. As explained in Remark 1, although mean estimation does not affect the convergence rate of our MCLE, Hosking (1996) shows that in the long memory setting it induces a negative $O(n^{-\beta})$ bias in the acf. The intuition is that since process is persistent, it oscillates around the sample average, even if it far away from the mean. This is interpreted as shorter memory. The lower is $\beta$, the stronger is this effect. We should note that while the bias is in principle also present for the fOU process, it is not visible. This is probably because the autoregressive parameter is increased in parallel with the Hurst index, forcing the process back toward its mean despite the increased dependence. In the Cauchy class, we do not exert direct control over the degree of mean reversion.
@@
In the close, we should point out that if one is concerned about the slow convergence rate of the mean estimator under long memory it is possible to transform the problem by first-differencing the data. While this drops information about the mean, it should reduce the bias in the estimation of the autocovariance-related parameters at the expense of an increased variance. However, as the simulation results suggest that the MCLE estimator performs adequately even with the original data in levels, also when the mean is estimated, we do not pursue this option here.

### v2

@@ 3.2 The Cauchy class
$\rho_{h}=(1+|h|^{2\alpha+1})^{-\beta/(2\alpha+1)},$ (3.6)
@@
Motivated by its ability to decouple the short- and long-term persistence, this process was studied as an empirical model for log-spot variance in Bennedsen, Lunde, and Pakkanen (2022). They fitted it with method-of-moments to high-frequency data from the E-mini S&P 500 futures contract and indeed found evidence of both roughness and long memory.
4 Monte Carlo analysis
4.1 Simulation design
@@
The fOU model was studied in Bolko, Christensen, Pakkanen, and Veliyev (2023) and Wang, Xiao, and Yu (2023), among others, to describe the stochastic log-variance of financial asset returns, $Y_{t}=\log\sigma_{t}^{2}$. [^15] Wang, Xiao, Yu, and Zhang (2025) prove the asymptotic theory for the time-domain MLE of the fOU process. In contrast to MCLE, the limiting distribution of the covariance-related part of $\hat{\theta}_{\text{MLE}}$ is normal with an $n^{-1/2}$ rate of convergence for every $H\in(0,1)$ whether or not the mean is estimated. The mean estimator itself has a slower convergence rate (continuous in $H$) for $H>1/2$, but it stays Gaussian. We follow this analogue here and adopt the Monte Carlo design of the former article by examining five different settings for $\theta_{\text{OU}}$ (listed in Table 1) ranging from very rough and short memory ($\alpha=-0.45$ or $H=0.05$) to smooth and long memory ($\alpha=0.20$ or $H=0.70$).
We include an equal number of distinct parameter values for $\theta_{\text{CC}}$ (listed in Table 3), keeping the roughness index fixed at the values of the Hurst exponent from the fOU process, i.e. $\alpha=-0.45,-0.40,\dots,0.20$, while moving the long-run persistence in the opposite direction, i.e. $\beta=0.25,0.50,\dots,1.25$. This design ensures that the latter features multiple settings with both roughness and long memory.
We simulate over the interval $[0,T]$, where $T$ can be interpreted as the number of days, setting $T=1095,1825,2555$. This amounts to three, five, and seven years worth of data. The largest value roughly matches the sample size in our empirical application to realized variance. We assume the process is observed discretely on an equispaced partition of $N$ points per day, so the $n=N\times T$ data vector is $(Y_{i\Delta})_{i=1,\dots,N\times T}$, where $\Delta=1/N$. We follow Christensen, Thyrsgaard, and Veliyev (2019) and base the analysis on $N=12$ observations per unit interval, yielding a new observation of the log-spot variance every second hour.
We simulate the Cauchy class via circulant embedding, which is an exact discretization that runs in $O(n\log n)$ time. However, since the ACF of the fOU needs to be evaluated with the help of numerical integration we opt for a slightly different approach for that model. In particular, at a generic time $t$ the solution of the fOU process in (3.4) implies that:
@@
To implement MCLE, we need to decide the terms to include in (2.4) through the set $Q$. We experiment here with triwise CLE (i.e. $q^{j}=q=3$) with $K=5$, where the elements of $k^{j}$ are separated by a fixed amount, i.e. $k^{j}=(0,\ell,2\ell)$ for $\ell=1,6,12,24,60$. This is a sufficient number of lags to guarantee the identification of our models. [^16] A routine, but tedious, derivation shows that the Cauchy class is identified with at least four different lags in the criterion function. The proof of this result is available at request. Identification of the fOU process is more difficult to establish, but it is examined in some length in Appendix B of Bolko, Christensen, Pakkanen, and Veliyev (2023). This setup means that $cl(\theta;y)$ is a summation of densities of the form $f(y_{i},y_{i+\ell},y_{i+2\ell};\theta)$. Apart from lag one with a two-hour gap, the other choices of $\ell$ correspond do a half-, whole-, two-, and five-day horizon. [^17] In a previous version of this article, we relied on a more “casual” implementation of CLE that incorporated various lags of the ACF, including ones at the very long end of the curve, based on “intuition.” However, upon a closer examination we noted that the performance of CLE deteriorates with such a configuration. We therefore switched to the current setup to boost the performance of CLE and better illuminate its promise. The reported results are rather robust to the exact configuration of the CL procedure. [^18] Appendix C reports results for the pairwise setting with $q=2$. The conclusion from this analysis is that triwise CLE systematically leads to lower RMSE in the rough setting, but the pairwise cousin can also be a superior alternative. The change of going from $q=2$ to $q=3$, or vice versa, is often small, however. We leave the pursuit of, arguably model-dependent, optimal choices of $q$ and $K$ to a future endeavor.
@@
We benchmark MCLE against a method-of-moments estimator (MME). We follow the implementation of Wang, Xiao, and Yu (2023) for the fOU process and Bennedsen, Lunde, and Pakkanen (2022) for the Cauchy class. [^19] We do not compare our MCLE approach to the GMM estimator of Bolko, Christensen, Pakkanen, and Veliyev (2023), since the latter relies on integrated variance, whereas we study spot variance. Neither do we benchmark it against the time-domain MLE of Wang, Xiao, Yu, and Zhang (2025) or the frequency-domain approximate likelihood of Fukasawa, Takabatake, and Westphal (2022) and Shi, Yu, and Zhang (2024) for the fOU, since our procedure cannot be expected to be superior to these in terms of efficiency. As we argue, MCLE trades (a slight) efficiency loss against (a much) faster runtime. The MME of Wang, Xiao, and Yu (2023), which is also extremely fast and has an asymptotic theory that closely resembles that of MCLE—with a discontinuity in the rate of convergence and the limiting distribution—represents a more level playing field and natural competitor in our view. The details are reviewed in Appendix B. It is worth pointing out that the first-step MME estimator of $\alpha$ (or $H$) relies on change-of-frequency approach based on second-differenced data. Hence, it does not depend on whether the mean is known or estimated. We should also emphasize that we start the MCLE routine with the MME as initial value. The number of Monte Carlo replications is 10,000.
4.2 Simulation outcome
The Monte Carlo evidence for the fOU process and Cauchy class are summarized separately in a series of tables below. In each one, we report the outcome of the MCLE next to that of the MME. For every setting, we show the bias of the parameter estimator (i.e., sample average less true value) and its standard error across replica (in parentheses). To compare MCLE with MME, we add the relative root mean squared error, i.e., $\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$. A number below (above) one indicates that MCLE is more (less) efficient than MME. [^20] The relative RMSE should be interpreted with caution. The actual RMSEs of MCLE and MME are often small, so it does not take much to move the ratio by a lot in either direction.
4.2.1 The fOU process
Looking at the results in Tables 12 for the fOU process first, we observe that both MCLE and MME perform well. The parameters are estimated with an immaterial bias and the standard errors are small, even with a limited sample. The sole exception is the mean-reversion parameter, $\kappa$, which is sometimes estimated with a slight upward bias, but it vanishes as the sample grows and is negligible for T=2, [^5] T=2{,}555. Whether or not the mean is estimated, the conclusion from Panels A – D is that MCLE is far superior to MME in terms of estimating the covariance-related parameters $(\kappa,\nu,H)$. This includes the relevant scenario with a Hurst exponent in the roughness territory. However, we also observe that its relative efficiency decreases as we move toward long memory. Indeed, in Panel E with $H=0.7$ the MCLE is slightly inferior to MME for $(\kappa,H)$ that determine the correlation structure of $Y$. This observation is consistent with the evidence for the Cauchy class. In contrast to that model, however, there is only a minuscule change in performance on $(\kappa,\nu,H)$ between the known and estimated $\mu$. In the latter setting, the estimation of $\mu$ itself is a dead draw.
Table 1: Parameter estimation of the fOU process [$q=3$ and $\mu$ known].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@ 4.2.1 The fOU process
$\mu$ 0.000
$\kappa$ 0.070 0.004 (0.018) 0.003 (0.016) 1.141 0.003 (0.014) 0.002 (0.012) 1.132 0.002 (0.012) 0.001 (0.010) 1.134
$\nu$ 0.200 0.001 (0.007) 0.000 (0.008) 0.927 0.000 (0.006) -0.000 (0.006) 0.921 0.000 (0.005) -0.000 (0.005) 0.913
$\alpha$ 0.200 0.001 (0.013) -0.000 (0.011) 1.229 0.000 (0.010) -0.000 (0.008) 1.230 0.000 (0.009) -0.000 (0.007) 1.224
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
Table 2: Parameter estimation of the fOU process [$q=3$ and $\mu$ estimated].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@
$\mu$ 0.000 0.000 (0.069) 0.000 (0.069) 1.001 0.000 (0.059) 0.000 (0.059) 1.001 0.000 (0.054) 0.000 (0.054) 1.000
$\kappa$ 0.070 0.012 (0.018) 0.009 (0.016) 1.157 0.008 (0.013) 0.006 (0.012) 1.148 0.007 (0.011) 0.005 (0.010) 1.151
$\nu$ 0.200 0.002 (0.007) 0.000 (0.008) 0.951 0.002 (0.006) -0.000 (0.006) 0.938 0.001 (0.005) -0.000 (0.005) 0.926
$\alpha$ 0.200 0.004 (0.013) -0.000 (0.011) 1.219 0.003 (0.010) -0.000 (0.008) 1.222 0.002 (0.009) -0.000 (0.007) 1.216
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
4.2.2 The Cauchy class
Table 3: Parameter estimation of the Cauchy class [$q=3$ and $\mu$ known].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@ 4.2.2 The Cauchy class
$\mu$ 0.000
$\beta$ 1.250 0.009 (0.109) 0.038 (0.184) 0.590 0.005 (0.085) 0.029 (0.146) 0.577 0.003 (0.072) 0.025 (0.125) 0.572
$\nu$ 0.200 -0.000 (0.006) -0.000 (0.006) 1.003 -0.000 (0.005) -0.000 (0.005) 1.003 -0.000 (0.004) -0.000 (0.004) 1.003
$\alpha$ 0.200 0.000 (0.009) 0.041 (0.011) 0.279 -0.000 (0.007) 0.041 (0.008) 0.222 -0.000 (0.006) 0.041 (0.007) 0.190
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
Table 4: Parameter estimation of the Cauchy class [$q=3$ and $\mu$ estimated].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
@@
Turning next to the Cauchy class, the conclusion is two-fold. First, we see from Table 3, where the mean is known, that the distribution of the covariance-related parameter estimators is again centered around the true value with minimal sampling variation for MCLE. The main difference is that we observe a slight upward bias in $\hat{\beta}_{\text{MCLE}}$, although it is well within a Monte Carlo standard error. By contrast, $\hat{\beta}_{\text{MME}}$ is regularly off target with a negative bias in Panel A reversing into a positive one in Panel E. It even moves in the wrong direction (toward a too low value) as $T$ increases for $\alpha=-0.45$ or $\alpha=-0.40$. The origin of this problem appears to be the first-stage MME of $\alpha$. As it turns out, in our numerical experiments $\hat{\alpha}_{\text{MME}}$ is biased for a fixed $\Delta$, and this causes the estimated persistence to deviate from the true one. In turn, this distorts the second-stage MME of $\beta$ with the latter acting as a suspension that pulls the model-implied ACF in the opposite direction. This explains the positive correlation between $\hat{\beta}_{\text{MME}}$ and $\hat{\alpha}_{\text{MME}}$. Since $\hat{\alpha}_{\text{MME}}$ is consistent in the infill asymptotic limit $\Delta\rightarrow 0$, this effect ought to disappear with more frequent sampling of the log-variance process, but that is of course not possible to enforce here, where $\Delta$ is fixed. This highlights a potential drawback of the two-step MME procedure. [^21] To investigate this issue further, in Appendix C we also report on a Monte Carlo study, where the volatility process is sampled once per day, i.e. $N=1$. The broad conclusion is that the bias problems in $\alpha$ with MME are intensified, indicating that it may be more imprecise with less frequently sampled data. This is especially troublesome, since the MME is often applied to daily RV (e.g. Gatheral, Jaisson, and Rosenbaum, 2018; Bennedsen, Lunde, and Pakkanen, 2022; Wang, Xiao, and Yu, 2023).
Second, as shown in Table 4, where the mean is estimated, $\hat{\beta}_{\text{MCLE}}$ exhibits a more pronounced upward bias than before, in particular for the extreme persistence configurations in Panels A – B, although it still dissipates as the sample size is increased. Measured by relative RMSE, the bias in $\hat{\beta}_{\text{MCLE}}$ is hurting a lot. Here, $\hat{\beta}_{\text{MME}}$ only has a slight bias, which is, however, again heading the wrong way as we collect more information. In Panel A, this is, in fact, initially helping to reduce the RMSE of $\hat{\beta}_{\text{MME}}$, because the bias crosses zero around T=1,825T=1{,}825. As consistent with the above, the sign of the bias in $\hat{\beta}_{\text{MME}}$ systematically is positive in Panels C – E and is, in fact, more pronounced than before. Here, $\hat{\beta}_{\text{MCLE}}$ is evidently a better estimator both in terms of bias and variance.
@@
The MCLE of the mean—a feasible Generalized Least Squares (GLS)-type weighted average—is unbiased across replica, but it exhibits a great deal of sampling variation and is typically way off target in each individual simulation, as evident from the magnitude of the standard error. [^22] The stationary Gaussian process in (2.1) is basically a linear regression on a constant, for which we impose a parametric assumption on the variance-covariance matrix of the disturbance term. As the mean is unrelated to the other parameters in the models we inspect, it follows from Magnus (1978) that the MLE of $\mu$ is the feasible GLS statistic calculated by plug-in of the MLEs of the parameters indexing $\Sigma$, i.e. $\hat{\mu}_{\text{MLE}}=(\iota^{\top}\hat{\Sigma}^{-1}\iota)^{-1}\iota^{\top}\hat{\Sigma}^{-1}Y_{n}^{\Delta}$, where $\iota=(1,\dots,1)^{\top}$. An inspection of the first-order condition $\partial cl(\theta;y)/\partial\mu=0$ reveals that, since each summand in the composite log-likelihood function is Gaussian, $\hat{\mu}_{\text{MCLE}}$ also has a GLS-type structure, but it entails a more complicated weighting scheme. The derivation of the exact expression is available at request. In essence, the mean estimator is very noisy, especially for persistent processes. As explained in Remark 1, although this does not affect the convergence rate, Hosking (1996) shows that in the long memory setting it induces a negative $O(n^{-\beta})$ bias in the ACF. The intuition is that when a process is highly persistent, even if it is far away from and never actually crosses the mean, it still tends to oscillate around the average. This is interpreted by the estimation procedure as shorter memory. The lower the $\beta$, the stronger is this effect. [^23] To what extent does the increased roughness in the Cauchy class of processes explain the deterioration in $\hat{\beta}_{\text{MCLE}}$ in the presence of long memory? To shed light on this, we performed a robustness check, i.e. an extra simulation with the most long memory, $\beta=0.25$ from Panel A, combined with the least roughness, $\alpha=0.20$ from Panel E, together with an intermediate standard deviation, $\nu=0.50$ from Panel C. In these experiments, the average $\hat{\beta}_{\text{MCLE}}$ value across different $T$ was around 0.37, so the bias was roughly cut in half relative to what we observe in Panel A of Table 4, but it remained elevated. We should note that while this effect is in principle also present for the fOU process, it is not visible. This is probably because the autoregressive parameter is increased in parallel with the Hurst index, forcing the process back toward its mean despite the increased dependence. In the Cauchy class, we do not exert direct control over the degree of mean reversion.
@@
In the close, we should point out that if one is concerned about the slow convergence rate of the mean estimator under long memory it is possible to transform the problem by first-differencing the data. While this drops information about the mean, it should reduce the bias in the estimation of the autocovariance-related parameters at the expense of an increased variance. However, as the simulation results suggest that the MCLE performs adequately even with the original data in levels, also when the mean is estimated, we do not pursue this option here.